x+3y+z=9
3x+y-z=-1
2x+2y-z=3

Answer by Alan3354(69443)

x+3y+z=9
3x+y-z=-1
2x+2y-z=3
------------
Add eqns 1 & 2, then eqns 1 & 3.
---> 2 eqns in x & y, no more z.
Go from there.